OpenStudy (anonymous):

sqrt 7 ( sqrt 14 + sqrt 3 )

OpenStudy (tkhunny):

What is to be done?

OpenStudy (anonymous):

sqrt 7 * sqrt 14 + sqrt 7 * sqrt 3 sqrt 98 + sqrt 21 7 sqrt 2 + sqrt 21

OpenStudy (tkhunny):

Just one thing, Why on Earth would you want to see \(\sqrt{98}\)? Keep things simpler! \(\sqrt{7}\cdot(\sqrt{14} + \sqrt{3}) = \sqrt{7}\cdot(\sqrt{7}\sqrt{2} + \sqrt{3}) = 7\sqrt{2} + \sqrt{21}\) No crazy numbers.

OpenStudy (anonymous):

Thank you I see what I was doing wrong

OpenStudy (tkhunny):

Well, it definitely wasn't "wrong". It was more of a style difference. Unique answers don't care how you find them. I just tend to think that we should not try to do things harder when they could be easier.

OpenStudy (anonymous):

I have one more question how do you get the sqrt symbol. I tried the equation box at the end of the post but it writes out sqrt.

OpenStudy (tkhunny):

I usually just type things myself. "The square root of seven" requires this (without the spaces or quotes) "\ ( \ s q r t { 7 } \ )"

OpenStudy (anonymous):

\(\sqrt{7}\) cool thank you

OpenStudy (tkhunny):

You can also right-click the math text, select "Show Math As" and pick "Tex Commands" to see the hiding coding secrets.